# TimeMosaic: Temporal Heterogeneity Guided Time Series Forecasting via Adaptive Granularity Patch and Segment-wise Decoding

### v2

@@ Appendix H Detailed Results
In accordance with prior works such as (Wu et al. 2023), we adopt the standard prediction lengths of $\{96,192,336,720\}$ for long-term forecasting tasks, and a fixed length of 12 for short-term forecasting.
@@
We further identify a common experimental flaw in many recent works: the test dataloader is often configured with drop_last=True, which can silently discard up to batch_size - 1 samples during evaluation. Although this issue has been highlighted in (Qiu et al. 2024), it remains present in recent literature (Chen et al. 2025; Wang et al. 2025a). To address this, we explicitly set drop_last=False in all our evaluations, ensuring complete coverage of the test set.
Dataset $d_{model}$ $d_{ff}$ #Layers #Heads Batch Size Epochs LR Patience
@@
  • Revin: Replaces z-score normalization with RevIN (Kim et al. 2021). We note that z-score normalization is also used for all experiments in Table 20.
@@
To further evaluate the scalability and generalization ability of TimeMosaic, we conduct large-scale experiments under the same settings as Table 21, comparing against representative Time Series Foundation Models (TSFMs). Specifically, TimeMosaic is trained with an input length of 512 and an output horizon of 720 on two NVIDIA V100 GPUs, requiring approximately 40 hours of training. The forecasting segment length is set to 16, with train_epochs=1, batch size of 256, a single Transformer layer, and 8 attention heads. The prompt embedding dimension is 8, and the adaptive patch search space is defined as [8, 16, 32, 64]. We adopt the BLAST dataset (Shao et al. 2025) for training. Under this configuration, the total number of trainable parameters is 27,259,924.
BLAST (BaLAnced Sampling Time series corpus) is a large-scale dataset specifically constructed for pre-training universal time series models. It consists of 321 billion observations, mainly from CMIP6 (Eyring et al. 2016) (32.5%), ERA5 (Hersbach et al. 2020) (30.0%), WeatherBench (Rasp et al. 2020) (25.7%), Buildings_900K (Emami, Sahu, and Graf 2023) (6.9%), and over 300 additional public datasets (Shao et al. 2025) (4.9%), covering diverse domains such as climate, meteorology, and energy. To ensure the independence of downstream evaluation, commonly used benchmark datasets (e.g., ETTh/ETTm, Weather, GlobalTemp) are deliberately excluded. In our experiments, we adopt random sampling on BLAST to generate training data, ensuring broad coverage of diverse patterns.

### v3

@@ Appendix H Detailed Results
In accordance with prior works such as (timesnet), we adopt the standard prediction lengths of $\{96,192,336,720\}$ for long-term forecasting tasks, and a fixed length of 12 for short-term forecasting.
@@
We further identify a common experimental flaw in many recent works: the test dataloader is often configured with drop_last=True, which can silently discard up to batch_size - 1 samples during evaluation. Although this issue has been highlighted in (qiu2024tfb), it remains present in recent literature (chen2025simpletm; wang2025fredf). To address this, we explicitly set drop_last=False in all our evaluations, ensuring complete coverage of the test set.
Dataset $d_{model}$ $d_{ff}$ Layers Heads Batch Size Epochs LR Patience
@@
  • Revin: Replaces z-score normalization with RevIN (kim2021reversible). We note that z-score normalization is also used for all experiments in Table 20.
@@
To further evaluate the scalability and generalization ability of TimeMosaic, we conduct large-scale experiments under the same settings as Table 21, comparing against representative Time Series Foundation Models (TSFMs). Specifically, TimeMosaic is trained with an input length of 512 and an output horizon of 720 on two NVIDIA V100 GPUs, requiring approximately 40 hours of training. The forecasting segment length is set to 16, with train_epochs=1, batch size of 256, a single Transformer layer, and 8 attention heads. The prompt embedding dimension is 8, and the adaptive patch search space is defined as [8, 16, 32, 64]. We adopt the BLAST dataset (blast) for training. Under this configuration, the total number of trainable parameters is 27,259,924.
BLAST (BaLAnced Sampling Time series corpus) is a large-scale dataset specifically constructed for pre-training universal time series models. It consists of 321 billion observations, mainly from CMIP6 (CMIP6) (32.5%), ERA5 (ERA5) (30.0%), WeatherBench (weatherbenchmark) (25.7%), Buildings_900K (emami2023buildingsbench) (6.9%), and over 300 additional public datasets (blast) (4.9%), covering diverse domains such as climate, meteorology, and energy. To ensure the independence of downstream evaluation, commonly used benchmark datasets (e.g., ETTh/ETTm, Weather, GlobalTemp) are deliberately excluded. In our experiments, we adopt random sampling on BLAST to generate training data, ensuring broad coverage of diverse patterns.